jafdskfe:

Laura Russo lost her earnings statement from Siler's Lawn and Garden. She recalls paying $43.00 in federal income taxes and knows she deposited $306.36 in her checking account. Each week she had deductions for Social Security and Medicare taxes: $22.35 for medical insurance; 1.5% state tax; 1.5% local tax; and $25.00 for the credit union. What is her gross weekly pay?

Razor:

306.36+ 43 = 349.36 <--- federal income tax 349.36 + 22.35 = 371.71 <--- medical insurance 371.71 + 25 = 396.71 <--- credit union 396.71 increased by 1.5% is 396.71 * (1.015) = 402.66 <--- state tax 402.66 increased by 1.5% is 402.66 * (1.015)= 409.13 <-- local tax. There may be 1 penny round-off error

jafdskfe:

so would 409.13 be the gross?

Razor:

Iiiiii think so...

jafdskfe:

okay thanks

Razor:

Welcome, close post when ready